BIMA ARRAY 3 MM SPECTRAL LINE SURVEY OF ORION-KL

\underline{D. N. FRIEDEL}, L. E. SNYDER, University of Illinois, Dept. of Astronomy, 1002 W. Green St, Urbana, IL 61801; B. E. TURNER, National Radio Astronomy Observatory, Charlottesville, VA 22903; ANTHONY J. REMIJAN, NASA Goddard Space Flight Center, Space and Earth Data Computing Division, Code 930, Greenbelt, MD 20771; National Research Council Resident Research Associate.

With the Berkeley-Illinois-Maryland-Association (BIMA) Array we have carried out a 3 mm (80-115.9 GHz) spectral line survey of Orion-KL. Orion-KL is composed of several distinct regions separated by only a few arcseconds, most notably the “Hot Core”, where N bearing species are most prevalent, and the “Compact Ridge”, where O bearing species are most prevalent. For our observations the average synthesized beam was 14”x8”. We detected over 700 spectral features from the “Hot Core” and over 600 spectral features from the “Compact Ridge”. We will present the abundances and spatial distribution of the detected species.